OpenStudy (anonymous):

A fair coin is tossed twice. Define the events. A = Getting a head on the first toss. B = Getting a head on the second toss. A and B = Getting a head on both the first and second tosses. A or B = Getting a head on the first toss, or the second toss, or both tosses. (a) Find P(A) = the probability of A. (b) Find P(B) = the probability of B. (c) Using the multiplication rule, find P(A and B) (d) Using the addition rule, find P(A or B). I got: a) 1/2 b) 1/2 c) 1/4 d) 1/8 Can someone please correct my answers and give me explanations on the ones I got wrong? Thank you!

OpenStudy (anonymous):

Your first three answers are right. I prefer to solve the last by saying that this outcome means one minus getting tails both times = 1-1/4 = 3/4 That's the answer, you need to use "addition rule."

OpenStudy (anonymous):

Thank you very much! I appreciate your help :)